OpenStudy (anonymous):

Matt and Emily are building a rectangular vegetable garden for their grandmother. They can use up to 60 feet of fencing to enclose the garden. Which of the following is a possible combination of length and width they can use to build the garden? A. Length: 5 feet, Width: 26 feet B. Length: 11 feet, Width: 20 feet C. Length: 12 feet, Width: 17 feet D. Length: 17 feet, Width: 14 feet

OpenStudy (anonymous):

2L + 2W = 60 plug in L and W and see if it holds true example A: L = 5 W = 26 2(5) + 2(26) = 62. 62 > 60 so A is NOT the answer. try the rest ^_^

OpenStudy (anonymous):

actually it is an inequality \[2L + 2W \le 60\] answer might be less than 60 and it would be right

OpenStudy (anonymous):

yeah idk it

OpenStudy (anonymous):

its plug and play.. plug in the given L and W and see if it is less than 60

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

all i get is 31 and one 21

OpenStudy (anonymous):

oops 1 29

OpenStudy (anonymous):

I think it is C because if you plug them in it = 58 and all the others when you plug them in = 62

OpenStudy (anonymous):

@musa19965 are you there?

OpenStudy (anonymous):

yeah i think of that as well i got a different answer

OpenStudy (anonymous):

only C works

OpenStudy (anonymous):

Yeah